# Decoding Geometric Properties in Non-Random Data from First Information-Theoretic Principles

### v2

@@ A.5.1 Algorithmic information dynamics in communication problems
Our theoretical results in Section A.3 show that the method employed to achieve the empirical findings in Section A.4 is sound with regard to the infinite asymptotic limit when computational resources are unbounded; and in addition it also shows the method can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Thus, our empirical results suggest that AID is capable of harnessing (under feasible computational resources) some of the algorithmic-informational properties that we demonstrated in ideal and generalised scenarios. As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, and most prominently for high algorithmically probable objects.
As explained in Section A.1, the (algorithmic) perturbations that we investigate in this work can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimensions’ lengths of a given partition to achieve another partition) independently, or can change both at the same time. The reader is invited to notice that the present work applies AID to the problem of communicating messages by performing a perturbation analysis on the received signal. Thereafter, the principles of AIT are applied through approximations to algorithmic complexity by using BDM. These approximations are employed as a performance measure that characterizes the criterion to estimate the correct multidimensional space by distinguishing it from the wrong ones, which are the outcomes of the perturbations. Therefore, as also raised in Section A.2, our present approach, or AID in general, cannot be reducible to a straightforward application of AIT (e.g. by using the algorithmic coding theorem to approximate the optimal distribution or another method to approximate algorithmic complexity [19]) so as to estimate the irreducible information content of the received signal in order to directly retrieve the original message sent by the emitter agent.
@@
In fact, it is the semi-computability that evinces the divergence power brought about by the landscape of the approximate complexity values, landscape which results from the perturbations and from which we can find the best candidates to choose. One could empirically estimate the values of the algorithmic complexities of each possible message and also compare them with each other in order to check which one displays the lowest algorithmic complexity, therefore more likely to be the correct one. However, due to the uncomputability of algorithmic complexity (or algorithmic probability), this approach alone will never assure that the estimation error is not necessarily high enough so as to make the complexity value of a message actually be lower than the complexity value found in another message, whose approximate complexity value initially displayed a lower value in comparison to that of the first message in the empirical estimation process. In other words, the uncomputable errors in any estimation of algorithmic complexity can deceive one into thinking a message which actually has higher complexity than another message, has a lower complexity than the latter. The objects in other dimensions may not actually be more random, they only appear more random to the approximating methods, but it is the perturbation analysis that helps us to find the intended patterns.
@@
Although it may seem paradoxical at first glance—because one may initially think that the errors produced by the approximating methods to the algorithmic complexity would be inherited by the perturbation analysis phase itself—, it is the perturbation analysis that enables one to overcome such a semi-computability limitation but only when looking at the whole landscape of complexity values one estimates with a sufficiently large amount of perturbations effected on the received signal. Enabled by the employment of AID in the problem of interpreting a signal without prior knowledge of its source, we believe this also is one of the contributions of our results. As the number of perturbed messages strongly dominate the number of possible best message candidates, our theoretical and empirical results show that perturbation analysis enables the downward spikes displayed in the landscape of complexities to more likely indicate the best message candidates, even though the exact values of the algorithmic complexities are uncomputable and one can only asymptotically approximate them from above.
@@
A.5.2 Toward an algorithmic noisy-channel coding theorem
The reader is invited to notice that Theorem 11 demonstrates an algorithmic-informational version of the noisy-channel coding theorem in classical information theory [20]. Instead of the purely stochastic noise to which the communication channel may be subject in the classical version, the communication problem we studied in this paper deals with communication channels under algorithmic noise, i.e., the case in which signals are prone (or may be subject) to algorithmic perturbations [26] that can change the originally sent signals like a computer program could do, then resulting in new signals—which may substantially differ from the respective original signals. This new version of the problem of communicating a message can be understood as being more general than the classical version because: it holds for zero-knowledge one-way555 Notice that the classical version also applies to one-way communication but since both agents must know each other’s encoding-decoding scheme beforehand, one may argue that at least some information was previously communicated between the emitter agent and the receiver agent, which by definition violates the condition of one-way communication in the first place. communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26].
@@ A.5.2 Toward an algorithmic noisy-channel coding theorem
A.5.3 Zero-knowledge communication and ZKPs
As shown in the main article and discussed in the previous Section A.5.2, one of the salient advantages of our results in Theorem 11 is the absence of a condition stating that the encoding-decoding scheme arbitrarily chosen by the emitter agent must be also known by the receiver agent. Such a condition that relaxes the need of prior knowledge is what defines the zero-knowledge characteristic of the problems we investigated in this work. Our results are not only agnostic to prior knowledge of encoding-decoding schemes used to encode the message by the emitter agent, but also to: the computation model and the programming language chosen in order to implement our proposed method; the computable (or semi-computable) method of approximation to algorithmic complexity; any arbitrarily chosen (computable) probability measure of the events; and the formal theory able to formalize the mathematical statements. We have demonstrated sufficient conditions in this zero-knowledge scenario for enabling the reconstruction of the original message (i.e., the original object embedded into the original multidimensional space to be conveyed to the receiver) in one-way communication channels.
In this manner, zero-knowledge communication (ZKC) occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes (and programming languages and computation models), and consequentially also to the original multidimensional space and object that are yet unknown by the receiver agent before any communication takes place. However, this zero-knowledge condition does not mean other assumptions regarding the unknown emitter are not being considered by the receiver (in addition to those mathematical conditions already investigated in Section A.3). For example, as it is the case studied in the present work, the assumption that the emitter agent is somehow capable of performing an arbitrary encoding and compression of the message into a signal stream.
@@ A.5.3 Zero-knowledge communication and ZKPs
In the ZKC case instead, one looks for high enough probabilities of one agent being correct about the very content of what the other agent intended to communicate, given that no information was known by the former about the latter. Intuitively, one agent (i.e., the receiver agent) should statistically guarantee that its own interpretation of the signal sent by the other agent (i.e., the emitter agent) to the former corresponds to the original message that the latter intended, and this understanding (i.e., decoding) should happen in such a way that no prior information (i.e., knowledge) about what and how the second agent will communicate in the first place is known.
@@
A.5.4 A perturbation analysis-based method for the maximum algorithmic likelihood
@@
In order to tackle the new challenges presented in Section A.2, we have investigated additional conditions that enable one to infer the correct message, given that the receiver has no prior knowledge about the source (i.e., object and multidimensional space). In this sense, one aspect of the contributions of the present work resides in demonstrating the existence of sufficient conditions that allows the application of the algorithmic coding theorem [39, 41] to estimate the best model or parameters, even in scenarios where algorithmic information distortions may take place due to isomorphic transformations of the multidimensional space. Being one of the central results and powerful tools from algorithmic information theory, the usage of the universal distribution established by the algorithmic coding theorem underpins methods based on universal (Solomonoff) induction or minimum description length  [27, 40], whose general solutions encompass and refine the maximum likelihood or maximum entropy methods [20, 40, 29].
@@
As discussed in Section A.5.1, another aspect of our empirical and theoretical results resides in showing how perturbation analysis can be employed to overcome some of the semi-computability limitations in the approximations to algoritmic complexity (or algorithmic probability) values. Because the algorithmic complexities of the perturbations dominate those of the best candidates, one can still statistically infer the most likely messages (associated with higher algorithmic probabilities) due to the increasing divergence between the actual complexity values of these best candidates and the perturbed signals. By taking the algorithmic complexity of the object and the multidimensional space as the parameters to be estimated, our methods and results demonstrate how one can statistically infer the parameters with highest algorithmic likelihood [33, 34] from perturbation analysis without the need of exactly calculating the (uncomputable) algorithmic complexity values. In this regard, as demonstrated in Theorem 11, this inference method is general enough to remain robust vis-a-vis any arbitrarily chosen alternative method to calculate the probabilities of the events, while assuming no bias in the priors (i.e, while taking a uniform distribution of the possible messages as starting point).
@@
For example, Theorem 11 shows that for any probability distribution, encoding-decoding scheme, programming language, and computation model, the algorithmic complexity of the correct message is strongly dominated by the algorithmic complexities of the perturbations, dominance which holds with (uniform) probability arbitrarily close to $1$. If one knows how to calculate an upper-bound approximation to the algorithmic complexities (approximations which can be calculated by computable methods), then there are sufficient conditions for which one can in general infer that the algorithmic complexity of the sent message increases much slower than the algorithmic complexity of the perturbations, or equivalently that the algorithmic probability of the sent message increases much faster than algorithmic complexity of the perturbed signals.
@@
Intuitively, as it is already known to hold analogously for Fisher information [20] (which gives a lower bound for the variance of any unbiased estimator) in traditional statistics, sharper and more isolated valleys (i.e., downward spikes) found in the landscape of complexity values calculated for each parameter should indicate a better approximation to the actual algorithmic complexity (or algorithmic probability) of the correct parameter. Formally, if the algorithmic complexities of the perturbed signals are known to be sufficiently higher than the approximations to the complexities of the best message candidates (along with the fact that the quantity of distinct perturbations also strongly dominates the quantity of best candidates), then Theorem 11 shows that the errors in these approximations are strongly dominated by (i.e., increase much slower than) the highest algorithmic complexity of the perturbations. Thus, finding the existence of more tighter bounds in the relationship between the variance of the algorithmic complexity change rate (e.g., a rate obtained by calculating the ratio between the information difference [37] and the algorithmic complexity of the perturbation that changed the elements of the system) and how small those errors are is a problem that demands future investigation. We believe that our present work also contributes by establishing first results that posit a new path of future research in this direction.

### v3

@@ Appendix B Algorithmic information dynamics in communication problems
Our theoretical results in Section III-B show that the method employed to achieve the empirical findings in Section II and in [30], and discussed in Section II-A, is sound with regard to the infinite asymptotic limit when computational resources are unbounded; and in addition it also shows the method can be generalised to multidimensional spaces in addition to bi- or tridimensional ones. Thus, our empirical results suggest that AID is capable of harnessing (under feasible computational resources) some of the algorithmic-informational properties that we demonstrated in ideal and generalised scenarios. As shown in [24, 40], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM, probability values which remain stable across distinct models of computation, and most prominently for high algorithmically probable objects.
@@
As fully explained in the Sup. Mat. (Section $1$) and in Section III, the (algorithmic) perturbations that we investigate in this work can change either the object (e.g., by flipping bits) or its multidimensional space (e.g., by reconfiguring the dimensions’ lengths of a given partition to achieve another partition) independently, or can change both at the same time. The reader is invited to notice that the present work applies AID to the problem of communicating messages by performing a perturbation analysis on the received signal. Thereafter, the principles of AIT are applied through approximations to algorithmic complexity by using BDM. These approximations are employed as a performance measure that characterizes the criterion to estimate the correct multidimensional space by distinguishing it from the wrong ones, which are the outcomes of the perturbations. Therefore, as also raised in Appendix A, our present approach, or AID in general, cannot be reducible to a straightforward application of AIT (e.g. by using the algorithmic coding theorem to approximate the optimal distribution or another method to approximate algorithmic complexity [19]) so as to estimate the irreducible information content of the received signal in order to directly retrieve the original message sent by the emitter agent. This is because it is in fact the semi-computability that evinces the divergence power brought about by the landscape of the approximate complexity values, landscape which results from the perturbations and from which we can find the best candidates to choose. One could empirically estimate the values of the algorithmic complexities of each possible message and also compare them with each other in order to check which one displays the lowest algorithmic complexity, therefore more likely to be the correct one. However, due to the uncomputability of algorithmic complexity (or algorithmic probability), this approach alone will never assure that the estimation error is not necessarily high enough to make the complexity value of a message actually be lower than the complexity value found in another message, whose approximate complexity value initially displayed a lower value in comparison to that of the first message in the empirical estimation process. In other words, the uncomputable errors in any estimation of algorithmic complexity can deceive one into thinking a message which actually has higher complexity than another message, has a lower complexity than the latter. The objects in other dimensions may not actually be more random, they only appear more random to the approximating methods, but it is the perturbation analysis that helps us to find the intended patterns.
@@
Although it may seem paradoxical at first glance—because one may initially think that all the errors produced by the approximating methods to the algorithmic complexity would be inherited by the perturbation analysis phase itself—, it is the perturbation analysis that enables one to overcome such a semi-computability limitation but only when looking at the whole landscape of complexity values one estimates with a sufficiently large amount of perturbations effected on the received signal. Enabled by the employment of AID in the problem of interpreting a signal without prior knowledge of its source, we believe this also is one of the contributions of our results. As the number of perturbed messages strongly dominate the number of possible best message candidates, our theoretical and empirical results show that perturbation analysis enables the downward spikes displayed in the landscape of complexities to more likely indicate the best message candidates, even though the exact values of the algorithmic complexities are uncomputable and one can only asymptotically approximate them from above.
@@
Appendix C Toward an algorithmic noisy-channel coding theorem
The reader is invited to notice that Theorem 1 demonstrates an algorithmic-informational version of the noisy-channel coding theorem in classical information theory [20]. Instead of the purely stochastic noise to which the communication channel may be subject in the classical version, the communication problem we studied in this paper deals with communication channels under algorithmic noise, i.e., the case in which signals are prone (or may be subject) to algorithmic perturbations [26] that can change the originally sent signals like a computer program could do, then resulting in new signals—which may substantially differ from the respective original signals. This new version of the problem of communicating a message can be understood as being more general than the classical version because: it holds for zero-knowledge one-way111 Notice that the classical version also applies to one-way communication but since both agents must know each other’s encoding-decoding scheme beforehand, one may argue that at least some information was previously communicated between the emitter agent and the receiver agent, which by definition violates the condition of one-way communication in the first place. communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26].
@@ Appendix C Toward an algorithmic noisy-channel coding theorem
Appendix D Zero-knowledge communication and ZKPs
As discussed in Appendix C, one of the salient advantages of our results in Theorem 1 is the absence of a condition stating that the encoding-decoding scheme arbitrarily chosen by the emitter agent must be also known by the receiver agent. Such a condition that relaxes the need of prior knowledge is what defines the zero-knowledge characteristic of the problems we investigated in this work. Our results are not only agnostic to prior knowledge of encoding-decoding schemes used to encode the message by the emitter agent, but also to: the computation model and the programming language chosen in order to implement our proposed method; the computable (or semi-computable) method of approximation to algorithmic complexity; any arbitrarily chosen (computable) probability measure of the events; and the formal theory able to formalize the mathematical statements. We have demonstrated sufficient conditions in this zero-knowledge scenario for enabling the reconstruction of the original message (i.e., the original object embedded into the original multidimensional space to be conveyed to the receiver) in one-way communication channels.
In this manner, zero-knowledge communication (ZKC) occurs when the receiver agent is able to correctly interpret the received signal as the originally intended message sent by the emitter agent given that the receiver has no knowledge about the encoding-decoding scheme chosen by the emitter. Notice that this condition of no prior information about the emitter agent only applies to encoding-decoding schemes (and programming languages and computation models), and consequentially also to the original multidimensional space and object that are yet unknown by the receiver agent before any communication takes place. However, this zero-knowledge condition does not mean other assumptions regarding the unknown emitter are not being considered by the receiver (in addition to those mathematical conditions already investigated in Section III-B). For example, as it is the case studied in the present work, the assumption that the emitter agent is somehow capable of performing an arbitrary encoding and compression of the message into a signal stream. A necessary assumption is that the signal sent to the receiver actually contains a message the emitter agent is trying to convey. To also ensure the conditions investigated enable our theoretical results to hold is that the receiver agent also needs to assume the emitter agent is not in fact trying to deceive the other end of the communication channel [44], as mentioned in Appendix A.
@@ Appendix D Zero-knowledge communication and ZKPs
In the ZKC case instead, one looks for high enough probabilities of one agent being correct about the very content of what the other agent intended to communicate, given that no information was known by the former about the latter. Intuitively, one agent (i.e., the receiver agent) should statistically (algorithmically or universally) guarantee that its own interpretation of the signal sent by the other agent (i.e., the emitter agent) to the former corresponds to the original message that the latter intended, and this understanding (i.e., decoding) should happen in such a way that no prior information (i.e., knowledge) about what and how the second agent will communicate in the first place is known.
@@
Appendix E A perturbation analysis-based method for the maximum algorithmic likelihood
@@
In order to tackle the new challenges presented in Appendix A, we have investigated additional conditions that enable one to infer the correct message, given that the receiver has no prior knowledge about the source (i.e., object and multidimensional space). In this sense, one aspect of the contributions of the present work resides in demonstrating the existence of sufficient conditions that allows the application of the algorithmic coding theorem [37, 39] to estimate the best model or parameters, even in scenarios where algorithmic information distortions may take place due to isomorphic transformations of the multidimensional space. Being one of the central results and powerful tools from algorithmic information theory, the usage of the universal distribution established by the algorithmic coding theorem underpins methods based on universal (Solomonoff) induction or minimum description length [27, 38], whose general solutions encompass and refine the maximum likelihood or maximum entropy methods [20, 38, 29].
@@
As discussed in Appendix B, another aspect of our empirical and theoretical results resides in showing how perturbation analysis can be employed to overcome some of the semi-computability limitations in the approximations to algoritmic complexity (or algorithmic probability) values. Because the algorithmic complexities of the perturbations dominate those of the best candidates, one can still statistically infer the most likely messages (associated with higher algorithmic probabilities) due to the increasing divergence between the actual complexity values of these best candidates and the perturbed signals. By taking the algorithmic complexity of the object and the multidimensional space as the parameters to be estimated, our methods and results demonstrate how one can statistically infer the parameters with highest algorithmic likelihood [41, 42] from perturbation analysis without the need of exactly calculating the (uncomputable) algorithmic complexity values. In this regard, as demonstrated in Theorem 1, this inference method is general enough to remain robust vis-a-vis any arbitrarily chosen alternative method to calculate the probabilities of the events, while assuming no bias in the priors (i.e, while taking a uniform distribution of the possible messages as starting point).
@@
For example, Theorem 1 shows that for any probability distribution, encoding-decoding scheme, programming language, and computation model, the algorithmic complexity of the correct message is strongly dominated by the algorithmic complexities of the perturbations, dominance which holds with (uniform) probability arbitrarily close to $1$. If one knows how to calculate an upper-bound approximation to the algorithmic complexities (approximations which can be calculated by computable methods), then there are sufficient conditions for which one can in general infer that the algorithmic complexity of the sent message increases much slower than the algorithmic complexity of the perturbations, or equivalently that the algorithmic probability of the sent message increases much faster than algorithmic probability of the perturbed signals.
@@
Intuitively, as it is already known to hold analogously for Fisher information [20] (which gives a lower bound for the variance of any unbiased estimator) in traditional statistics, sharper and more isolated valleys (i.e., downward spikes) found in the landscape of complexity values calculated for each parameter should indicate a better approximation to the actual algorithmic complexity (or algorithmic probability) of the correct parameter. Formally, if the algorithmic complexities of the perturbed signals are known to be sufficiently higher than the approximations to the complexities of the best message candidates (along with the fact that the quantity of distinct perturbations also strongly dominates the quantity of best candidates), then Theorem 1 shows that the errors in these approximations are strongly dominated by (i.e., increase much slower than) the highest algorithmic complexity of the perturbations. Thus, finding the existence of more tighter bounds in the relationship between the variance of the algorithmic complexity change rate (e.g., a rate obtained by calculating the ratio between the information difference [33] and the algorithmic complexity of the perturbation that changed the elements of the system) and how small those errors are is a problem that demands future investigation. In contrast to traditional universal induction methods in AIT that looks for the optimal solution (or model) by directly maximising the algorithmic probability of the models (or minimising the complexity of the probability distribution that best fits the data), we believe that our present work also contributes by establishing first results that posit a new path of future research in this direction of investigating the global shape of the landscape of complexities/probabilities of alternative events or models.